# When is an Embedder More Promising than Another?

### v1

@@ 3 Quantifying Information Sufficiency Between Embedding Models
We want to compare embedding models using the concept of deficiency, leveraging Prop. 2 and Corollary 1. These propositions suggest that the performance on any classification task of an embedding model $U$ relative to the model $Z$ is bounded by $\delta(P_{U|X}\rightarrow P_{Z|X})$. However, estimating the deficiency from data samples is notably challenging [95], and while upper bounds derivation exists, they do not necessarily make it tractable.
@@ 3.1 Estimating Information Sufficiency
The deficiency $\delta(P_{U|X}\rightarrow P_{Z|X})$ between two embedding models $P_{U|X}$ and $P_{Z|X}$, measures how well $U$ can be used to simulate $Z$ using a Markov kernel $M\in\mathcal{K}(\mathsf{Z}|\mathsf{U})$. This section aims to build a tractable proxy for this reconstruction cost. To this end, we estimate how much we can reduce the uncertainty about $Z$ by observing $U$ by learning an appropriate Markov kernel. This corresponds to the information sufficiency [29, 4] and can be interpreted as the information-theoretic counterpart of the deficiency. The information deficiency between $U$ and $Z$ is then defined as:
@@
The information sufficiency $\mathcal{I}_{S}(U\rightarrow Z)$, relative to parametric classes of distributions $\mathcal{F}_{\Theta}(\mathsf{Z})$ and $\mathcal{K}_{\Theta}(\mathsf{Z}|\mathsf{U})$ (multivariate Gaussian mixtures [82]) is defined:
$\mathcal{I}_{S}(U\rightarrow Z)\triangleq\underbrace{\inf_{f\in\mathcal{F}_{\Theta}(\mathsf{Z})}\mathbb{E}\left[-\log f(Z)\right]}_{\text{Uncertainty of }Z}-\underbrace{\mathbb{E}\left[\inf_{M\in\mathcal{K}_{\Theta}(\mathsf{Z}|\mathsf{U})}\mathbb{E}\left[-\log M(Z|U)|U\right]\right]}_{\text{Uncertainty when simulating }Z\text{ from }U\text{ with }M}.$ (1)
@@
When the information sufficiency $\mathcal{I}_{S}(U\rightarrow Z)$ is large, it signifies that $U$ offers a substantial amount of information to simulate $Z$, a proxy for a small deficiency. Conversely, when $\mathcal{I}_{S}(U\rightarrow Z)$ is lower, it implies that the channel $P_{Z|Y}$ is subject to considerable noise or randomness, leading to a greater loss of statistical information.
We hence attempt to simulate $Z$ from $U$ by learning a Markov kernel $M\in\mathcal{K}_{\Theta}(\mathsf{Z}|\mathsf{U})$, via a mixture of multivariate Gaussians, and measure the uncertainty reduction it induces.
@@
For set of embbeders $(Z_{k})_{k}$ represented by their Markov kernels $\{P_{Z_{k}|X}\}_{k}$, we compute the pairwise information sufficiency $\mathcal{I}_{S}(Z_{k}\rightarrow Z_{l})$. The pairwise information sufficiency matrix defines the adjacency matrix of a directed graph of embedders (Figure 2). Corollary 1 shows that embedders sharing high information sufficiency are expected to perform similarly on any downstream tasks, motivating the identification of communities in the graph. While the graph construction is in $\mathcal{O}(N^{2})$; where $N$ is the number of embedders, it is in practice tractable for a reasonable number of embedders (the reader is referred to Sec. E.5) for more details).
@@
We construct the set of all information sufficiency using $Z_{k}$: $\mathcal{S}_{\mathcal{I_{S}}}\left(k\right)=\left\{\mathcal{I}_{S}(Z_{k}\rightarrow Z_{l})\right\}_{l\not=k}$. We build our information sufficiency score ($\overline{\mathcal{I}_{S}}$ score) by taking the median of $\mathcal{S}_{\mathcal{I_{S}}}\left(k\right)$. Details on the $\overline{\mathcal{I}_{S}}$ score’s estimation, such as the motivation behind the median choice, can be found in Sec. E.1.
@@ 4 Experimental Setup
    We compare the models’ performances on the downstream tasks and the $\overline{\mathcal{I}_{S}}$ score by measuring three types of correlations: the Pearson correlation, the Spearman correlation, and the Kendall-Tau coefficient. [^3] For the experiments in molecular modeling, in each subset regression and classification tasks are mixed. Hence, we do not compute the Pearson correlation to avoid mixing scores obtained for different metrics. [^2] The code used to perform all experiments is available at https://anonymous.4open.science/r/emir-B8D3/Readme.md
@@
(b) Molecular Modelling
Figure 3: Correlation between $\overline{\mathcal{I}_{S}}$ scores and downstream task performances in (a) NLP and (b) Molecular Modelling. $\varrho_{p}$ is the Pearson correlation, $\varrho_{s}$ is the spearman correlation, and $\tau$ is the Kendall-Tau coefficient. See Sec. C.3.1 for unaggregated results in NLP and Sec. D.3 in molecular modeling.
@@ Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011013290R2 made by GENCI, and enabled by support provided by Calcul Quebec and the Digital Research Alliance of Canada. We warmly thank Heitor Rapela and Banafsheh Karimian for their advice and comments about our work. We also owe a special highlight to Loïc Fosse for the many discussions and hindsights he provided and for the subsequent follow-up projects.
@@ Appendix A Background and Notation
We define the total variation distance between two Markov kernels $T,T^{\prime}\in\mathcal{K}(\mathsf{Y}|\mathsf{X})$ as follows:
$\|T-T^{\prime}\|_{\textrm{TV}}\triangleq\sup_{x\in\mathsf{X}}\|T(\cdot|x)-T^{\prime}(\cdot|x)\|_{\textrm{TV}}.$ (7)
@@ Appendix B Proofs Theoretical Results
Proposition 3 (Relationships of sufficiency and information).
@@
    Informativeness $\Rightarrow$ higher capacity to distinguish concepts. If the embedding model $P_{U|X}$ is more informative than embedding model $P_{Z|X}$, i.e. $U\succcurlyeq_{I}Z$, then
    $\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\geqslant\textrm{KL}\big{(}P_{Z|Y}(\cdot|y_{0})\|P_{Z|Y}(\cdot|y_{1})\big{)},$
@@ B.1 Proof Proposition 1
It is immediate to check that the data-processing inequality and the Markov chain $Y\leftrightarrow X\leftrightarrow U\leftrightarrow Z$ implies the relation in claim (i). On the other hand, (ii) is proved by means of an explicit counterexample [57]. Given any $20<p<1/$ 2 with $\bar{p}=1-p$. For some $\epsilon,\delta>0$, consider two discrete embedding models defined by the following matrices:
$P_{Z|X}=P_{U|X}\left(\begin{array}[]{cc}1+\epsilon&1-\epsilon\\ 1/2&1/2\end{array}\right)\quad\textrm{with}\quad P_{U|X}=\left(\begin{array}[]{cc}p&\bar{p}\\ p+\delta&\bar{p}-\delta\end{array}\right).$ (8)
By taking $\epsilon,\delta>0$ small enough, both $P_{U|X}$ and $P_{Z|X}$ are stochastic matrices. It follows that $P_{Z|X}$ is not a degraded version of $P_{U|X}$ but provided that $\epsilon,\delta$ are sufficient small, the embedding model $P_{U|X}$ is more informative than $P_{Z|X}$, which proves the claim.
In order to show (iii), let $P_{U|Y}$ and $P_{Z|Y}$ be the corresponding probability measures induced by $P_{X|U}(\cdot|u)$ via the embedding models:
@@
$P_{Z|Y}(Z|y)=\int_{\mathsf{X}}P_{Z|X}(Z|x)P_{X|Y}(\mathsf{d}x|y),\quad\forall y\in\mathsf{Y},\,Z\in\mathcal{B}(\mathsf{Z}),$ (10)
@@
$P_{Z|Y}(Z|y)=\mathbbm{1}[y=y_{0}]P_{Z|Y}(Z|y_{0})+\mathbbm{1}[y=y_{1}]P_{Z|Y}(Z|y_{1})$ (12)
and let $P_{Y}(y_{0})=\lambda$ and $P_{Y}(y_{1})=1-\lambda$. The above probability measures correspond to a quadruple of random variables: $(Y_{\lambda},X_{\lambda},U_{\lambda},Z_{\lambda})\in\mathcal{P}(\mathsf{Y}\times\mathsf{X}\times\mathsf{U}\times\mathsf{Z})$. Consider the function $f(\lambda)$ defined by
$f(\lambda)=I(Y_{\lambda};U_{\lambda})-I(Y_{\lambda};Z_{\lambda}).$
@@
$f^{\prime}(0)=\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}-\textrm{KL}\big{(}P_{Z|Y}(\cdot|y_{0})\|P_{Z|Y}(\cdot|y_{1})\big{)}\geqslant 0,$
which implies the claim (iii). This concludes the proof. ∎
@@ B.2 Comments about capacity to distinguish concepts
$\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\geqslant 2\|P_{Z|Y}(\cdot|y_{0})-P_{Z|Y}(\cdot|y_{1})\|_{\textrm{TV}}^{2}.$
From which, it is easy to verify that the accuracy of the expected Bayes accuracy of the optimal classifier based on $Z$ is upper bounded by [108, Lemma 2.1]:
$\sup_{\psi}\Pr(\psi(Z)=Y)\leqslant 1-\frac{1}{2}\exp\left(-\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\right),$
@@ B.3 Proof of Proposition 2 and Corollary 1
Clearly, the assumption (i) implies the statement (ii) by Data-Processing. Conversely, let us assume point (ii) holds. This means that, for all joint probability distribution $P_{YX}$, there exists $\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})$ such that
@@
$\quad\quad\quad\quad\quad\sup\limits_{\rho_{Z}}\displaystyle\sum_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\int_{\mathsf{Z}}\rho_{Z}(y|z)P_{Z|X}(\mathsf{d}z|x),$ (13)
where $\rho_{Z}:\mathsf{Z}\rightarrow\mathcal{P}(\mathsf{Y})$ is a (possibly randomized) inference procedure is transition probabilities, which the learner can optimize to maximize the guessing probability.
Let the decision rule $\rho_{Z}(y|z)=\mathbb{1}[z\in{A}_{y}]$ for any partition $\{{A}_{y}\}_{y\in\mathsf{Y}}$ of $\mathsf{Z}$ with ${A}_{y}\in\mathcal{B}(\mathsf{Z})$. Then, for any $P_{YX}$, expression (13) implies the existence there exists $\rho_{U}(y|u)$ such that
$\displaystyle\sum\limits_{(y,x)\in\mathsf{Y}\in\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{\mathsf{Z}}P_{Z|X}(\mathsf{d}z|x)\mathbb{1}[z\in{A}_{y}]-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]$ (14)
$=\quad\quad\displaystyle\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{{A}_{y}}P_{Z|X}(\mathsf{d}z|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]\leqslant 0.$ (15)
@@
$\sup_{P_{XY}}\inf_{\rho_{U}}\displaystyle\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{\mathcal{A}_{y}}P_{Z|X}(\mathsf{d}z|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]\leqslant 0.$ (16)
@@
$\inf_{\rho_{U}}\,\sup_{P_{YX}}\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\Gamma\big{(}(y,x),\rho_{U}\big{)}\leqslant 0$ (17)
$\Gamma\big{(}(y,x),\rho_{U}\big{)}\triangleq\left[\displaystyle\int_{\mathcal{A}_{y}}P_{Z|X}(\mathsf{d}z|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right].$ (18)
@@
$\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)}=0,$
and thus,
$\max\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)}\geqslant 0,$
since by contradiction otherwise
$\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)}<0.$
@@
$\inf\limits_{\rho_{U}}\sup\limits_{P_{YX}}\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{\mathcal{A}_{y}}P_{Z|X}(\mathsf{d}z|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]$
$=\inf\limits_{\rho_{U}}\sup\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)},$ (19)
which means the maximum is achieved by degenerate random variables $(Y,X)$ achieving the maximum. Consequently, we have that
$\inf_{\rho_{U}}\sup_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)}=0,$
which implies $\|P_{Z|X}-\rho_{U}\!\circ\!P_{U|X}\|_{\textrm{TV}}=0$ for all partition, and so the existence of the transition probability $\rho_{U}$ such that $P_{U|X}$ is sufficient for $P_{Z|X}$. This concludes the proof of the Proposition 2. ∎
@@
$\inf_{\rho_{U}}\sup_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)}\leqslant\varepsilon.$ (20)
@@
$\max_{x\in\mathsf{X}}\,\sum\limits_{y\in\mathsf{Y}}|\Gamma\big{(}(y,x),\rho_{U}\big{)}|.$
The above quantity is the induced $\ell_{1}$-norm distance between $\int_{\mathcal{A}_{y}}P_{Z|X}(\mathsf{d}z|x)$ and $\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)$. Since, for all $x\in\mathsf{X}$,
@@
$\sum\limits_{y\in\mathsf{Y}}|\Gamma\big{(}(y,x),\rho_{U}\big{)}|=2\sum\limits_{y\in\mathsf{Y}:\Gamma((y,x),\rho_{U})\geqslant 0}|\Gamma\big{(}(y,x),\rho_{U}\big{)}|,\quad\textrm{for all $x\in\mathsf{X}$}$
@@
$\sum\limits_{y\in\mathsf{Y}}|\Gamma\big{(}(y,x),\rho_{U}\big{)}|\leqslant 2|\mathsf{X}|\max\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}\Gamma\big{(}(y,x),\rho_{U}\big{)}\leqslant 2\varepsilon|\mathsf{X}|,\quad\textrm{for all $x\in\mathsf{X}$}$

### v2

@@ 3 Quantifying Information Sufficiency Between Embedding Models
We want to compare embedding models using the concept of deficiency, leveraging Prop. 2 and Corollary 1. These propositions suggest that the performance on any classification task of an embedding model $U$ relative to the model $V$ is bounded by $\delta(P_{U|X}\rightarrow P_{V|X})$. However, estimating the deficiency from data samples is notably challenging [95], and while upper bounds derivation exists, they do not necessarily make it tractable.
@@ 3.1 Estimating Information Sufficiency
The deficiency $\delta(P_{U|X}\rightarrow P_{V|X})$ between two embedding models $P_{U|X}$ and $P_{V|X}$, measures how well $U$ can be used to simulate $V$ using a Markov kernel $M\in\mathcal{K}(\mathsf{V}|\mathsf{U})$. This section aims to build a tractable proxy for this reconstruction cost. To this end, we estimate how much we can reduce the uncertainty about $Z$ by observing $U$ by learning an appropriate Markov kernel. This corresponds to the information sufficiency [29, 4] and can be interpreted as the information-theoretic counterpart of the deficiency. The information deficiency between $U$ and $V$ is then defined as:
@@
The information sufficiency $\mathcal{I}_{S}(U\rightarrow V)$, relative to parametric classes of distributions $\mathcal{F}_{\Theta}(\mathsf{V})$ and $\mathcal{K}_{\Theta}(\mathsf{V}|\mathsf{U})$ (multivariate Gaussian mixtures [82]) is defined:
$\mathcal{I}_{S}(U\rightarrow V)\triangleq\underbrace{\inf_{f\in\mathcal{F}_{\Theta}(\mathsf{V})}\mathbb{E}\left[-\log f(V)\right]}_{\text{Uncertainty of }V}-\underbrace{\mathbb{E}\left[\inf_{M\in\mathcal{K}_{\Theta}(\mathsf{V}|\mathsf{U})}\mathbb{E}\left[-\log M(V|U)|U\right]\right]}_{\text{Uncertainty when simulating }V\text{ from }U\text{ with }M}.$ (1)
@@
When the information sufficiency $\mathcal{I}_{S}(U\rightarrow V)$ is large, it signifies that $U$ offers a substantial amount of information to simulate $V$, a proxy for a small deficiency. Conversely, when $\mathcal{I}_{S}(U\rightarrow V)$ is lower, it implies that the channel $P_{V|Y}$ is subject to considerable noise or randomness, leading to a greater loss of statistical information.
We hence attempt to simulate $V$ from $U$ by learning a Markov kernel $M\in\mathcal{K}_{\Theta}(\mathsf{V}|\mathsf{U})$, via a mixture of multivariate Gaussians, and measure the uncertainty reduction it induces.
@@
For set of embbeders $(Z_{k})_{k}$ represented by their Markov kernels $\{P_{Z_{k}|X}\}_{k}$, we compute the pairwise information sufficiency $\mathcal{I}_{S}(Z_{k}\rightarrow Z_{l})$. The pairwise information sufficiency matrix defines the adjacency matrix of a directed graph of embedders (Figure 2). Corollary 1 shows that embedders sharing high information sufficiency are expected to perform similarly on any downstream tasks, motivating the identification of communities in the graph. While the graph construction is in $\mathcal{O}(N^{2})$; where $N$ is the number of embedders, it is in practice tractable for a reasonable number of embedders (refer to Sec. E.6) for more details).
@@
We construct the set of all information sufficiency using $Z_{k}$: $\mathcal{S}_{\mathcal{I_{S}}}\left(k\right)=\left\{\mathcal{I}_{S}(Z_{k}\rightarrow Z_{l})\right\}_{l\not=k}$. We build our information sufficiency score ($\overline{\mathcal{I}_{S}}$ score) by taking the median of $\mathcal{S}_{\mathcal{I_{S}}}\left(k\right)$. Details on the $\overline{\mathcal{I}_{S}}$ score’s estimation can be found in Sec. E.1.
@@ 4 Experimental Setup
    We compare the models’ performances on the downstream tasks and the $\overline{\mathcal{I}_{S}}$ score by measuring three types of correlations: the Pearson correlation, the Spearman correlation, and the Kendall-Tau coefficient. [^3] For the experiments in molecular modeling, in each subset regression and classification tasks are mixed. Hence, we do not compute the Pearson correlation to avoid mixing scores obtained for different metrics.(See  Sec. E.5 for additional baselines).
@@
(b) Molecular Modeling
Figure 3: Correlation between $\overline{\mathcal{I}_{S}}$ scores and downstream task performances in (a) NLP and (b) Molecular Modelling. $\varrho_{p}$ is the Pearson correlation, $\varrho_{s}$ the spearman correlation, and $\tau$ is the Kendall-Tau coefficient. See Sec. C.3.1 for unaggregated results in NLP and Sec. D.3 in molecular modeling.
@@ Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011013290R2 made by GENCI, and enabled by support provided by Calcul Quebec and the Digital Research Alliance of Canada. We warmly thank Heitor Rapela, Banafsheh Karimian, and Eric Aubinais for their advice and comments about our work. We also owe a special highlight to Loïc Fosse for the many discussions and hindsights he provided and for the subsequent follow-up projects.
@@ Appendix A Background and Notation
We define the average of the total variation distance between two Markov kernels $T,T^{\prime}\in\mathcal{K}(\mathsf{Y}|\mathsf{X})$ as follows:
$\mathbb{E}\|T-T^{\prime}\|_{\textrm{TV}}\triangleq\mathbb{E}\|T(\cdot|X)-T^{\prime}(\cdot|X)\|_{\textrm{TV}}.$ (7)
@@ Appendix B Proofs Theoretical Results
Proposition (Relationships of sufficiency and information).
@@
    Informativeness $\Rightarrow$ higher capacity to distinguish concepts. If the embedding model $P_{U|X}$ is more informative than embedding model $P_{Z|X}$, i.e. $U\succcurlyeq_{I}V$, then
    $\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\geqslant\textrm{KL}\big{(}P_{V|Y}(\cdot|y_{0})\|P_{V|Y}(\cdot|y_{1})\big{)},$
@@ B.1 Proof Proposition 1
It is immediate to check that the data-processing inequality and the Markov chain $Y\leftrightarrow X\leftrightarrow U\leftrightarrow V$ implies the relation in claim (i). On the other hand, the non-equivalence is proved by means of an explicit counterexample [57]. Given any $20<p<1/$ 2 with $\bar{p}=1-p$. For some $\epsilon,\delta>0$, consider two discrete embedding models defined by the following matrices:
$P_{V|X}=1/2\left(\begin{array}[]{cc}1+\epsilon&1-\epsilon\\ 1&1\end{array}\right)P_{U|X}\quad\textrm{with}\quad P_{U|X}=\left(\begin{array}[]{cc}p&\bar{p}\\ p+\delta&\bar{p}-\delta\end{array}\right).$ (8)
By taking $\epsilon,\delta>0$ small enough, both $P_{U|X}$ and $P_{V|X}$ are stochastic matrices. It follows that $P_{V|X}$ is not a degraded version of $P_{U|X}$ but provided that $\epsilon,\delta$ are sufficient small, the embedding model $P_{U|X}$ is more informative than $P_{V|X}$, which proves the claim.
In order to show (ii), let $P_{U|Y}$ and $P_{V|Y}$ be the corresponding probability measures induced by $P_{X|U}(\cdot|u)$ via the embedding models:
@@
$P_{V|Y}(V|y)=\int_{\mathsf{X}}P_{V|X}(V|x)P_{X|Y}(\mathsf{d}x|y),\quad\forall y\in\mathsf{Y},\,V\in\mathcal{B}(\mathsf{V}),$ (10)
@@
$P_{V|Y}(V|y)=\mathbbm{1}[y=y_{0}]P_{V|Y}(V|y_{0})+\mathbbm{1}[y=y_{1}]P_{V|Y}(V|y_{1})$ (12)
and let $P_{Y}(y_{0})=\lambda$ and $P_{Y}(y_{1})=1-\lambda$. The above probability measures correspond to a quadruple of random variables: $(Y_{\lambda},X_{\lambda},U_{\lambda},V_{\lambda})\in\mathcal{P}(\mathsf{Y}\times\mathsf{X}\times\mathsf{U}\times\mathsf{V})$. Consider the function $f(\lambda)$ defined by
$f(\lambda)=I(Y_{\lambda};U_{\lambda})-I(Y_{\lambda};V_{\lambda}).$
@@
$f^{\prime}(0)=\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}-\textrm{KL}\big{(}P_{V|Y}(\cdot|y_{0})\|P_{V|Y}(\cdot|y_{1})\big{)}\geqslant 0,$
which implies the claim (ii). This concludes the proof. ∎
@@ B.2 Comments about capacity to distinguish concepts
$\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\geqslant 2\|P_{V|Y}(\cdot|y_{0})-P_{V|Y}(\cdot|y_{1})\|_{\textrm{TV}}^{2}.$
From which, it is easy to verify that the accuracy of the expected Bayes accuracy of the optimal classifier based on $V$ is upper bounded by [108, Lemma 2.1]:
$\sup_{\psi}\Pr(\psi(V)=Y)\leqslant 1-\frac{1}{2}\exp\left(-\textrm{KL}\big{(}P_{U|Y}(\cdot|y_{0})\|P_{U|Y}(\cdot|y_{1})\big{)}\right),$
@@ B.3 Proof of Proposition 2 and Corollary 1
Clearly, the assumption (i) implies the statement (ii) by Data-Processing. Conversely, let us assume point (ii) holds. This means that, for every probbaility distribution $P_{X}$ and all conditional probability distributions $P_{Y|X}$, there exists $\rho_{U}:\mathsf{U}\rightarrow\mathcal{P}(\mathsf{Y})$ such that
@@
$\quad\quad\quad\quad\quad\sup\limits_{\rho_{V}}\displaystyle\sum_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\int_{\mathsf{V}}\rho_{V}(y|v)P_{V|X}(\mathsf{d}v|x),$ (13)
where $\rho_{V}:\mathsf{V}\rightarrow\mathcal{P}(\mathsf{Y})$ is a (possibly randomized) inference procedure is transition probabilities, which the learner can optimize to maximize the guessing probability.
Let the decision rule $\rho_{V}(y|v)=\mathbb{1}[v\in{A}_{y}]$ for any partition $\{{A}_{y}\}_{y\in\mathsf{Y}}$ of $\mathsf{V}$ with ${A}_{y}\in\mathcal{B}(\mathsf{V})$. Then, for any $P_{Y|X}$, expression (13) implies the existence there exists $\rho_{U}(y|u)$ such that
$\displaystyle\sum\limits_{(y,x)\in\mathsf{Y}\in\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{\mathsf{V}}P_{V|X}(\mathsf{d}v|x)\mathbb{1}[v\in{A}_{y}]-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]$ (14)
$=\quad\quad\displaystyle\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{{A}_{y}}P_{V|X}(\mathsf{d}v|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]\leqslant 0.$ (15)
@@
$\sup_{P_{Y|X}}\inf_{\rho_{U}}\displaystyle\sum\limits_{(y,x)\in\mathsf{Y}\times\mathsf{X}}P_{YX}(y,x)\left[\displaystyle\int_{\mathcal{A}_{y}}P_{V|X}(\mathsf{d}v|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right]\leqslant 0.$ (16)
@@
$\inf_{\rho_{U}}\sup_{\{A_{y}\}}\,\mathbb{E}\left[\sup_{P_{Y|X}}\sum\limits_{y\in\mathsf{Y}}P_{Y|X}(y|X)\Gamma\big{(}(y,X),\rho_{U}\big{)}\right]\leqslant 0$ (17)
$\Gamma\big{(}(y,x),\rho_{U}\big{)}\triangleq\left[\displaystyle\int_{\mathcal{A}_{y}}P_{V|X}(\mathsf{d}v|x)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)\right].$ (18)
@@
$\sum\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,x),\rho_{U}\big{)}=0,$ (19)
for each $x\in\mathsf{X}$ and thus,
$\max\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,x),\rho_{U}\big{)}\geqslant 0,$ (20)
where the equality holds if and only if $\Gamma\big{(}(y,x),\rho_{U}\big{)}=0$ for all $y\in\mathsf{Y}$, for each $x\in\mathsf{X}$, since by contradiction otherwise
$\sum\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,x),\rho_{U}\big{)}<0.$ (21)
@@
$\inf\limits_{\rho_{U}}\sup\limits_{\{A_{y}\}}\mathbb{E}\left[\sup\limits_{P_{Y|X}}\sum\limits_{y\in\mathsf{Y}}P_{Y|X}(y|X)\left(\displaystyle\int_{\mathcal{A}_{y}}P_{V|X}(\mathsf{d}v|X)-\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|X)\right)\right]$
$=\inf\limits_{\rho_{U}}\sup\limits_{\{A_{y}\}}\mathbb{E}\left[\max\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,X),\rho_{U}\big{)}\right],$ (22)
which means the maximum is achieved by degenerate random variables $Y=f(X)$ achieving the maximum for each $x\in\mathsf{X}$. Consequently, we have that
$\inf\limits_{\rho_{U}}\mathbb{E}\|P_{V|X}-\rho_{U}\!\circ\!P_{U|X}\|_{\textrm{TV}}$ $=$ $\inf\limits_{\rho_{U}}\sup\limits_{\{A_{y}\}}\,\mathbb{E}\left[\max\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,X),\rho_{U}\big{)}\right]=0,$ (23)
hence $\inf\limits_{\rho_{U}}\mathbb{E}\|P_{V|X}-\rho_{U}\!\circ\!P_{U|X}\|_{\textrm{TV}}=0$, and so the existence of the transition probability $\rho_{U}$ such that $P_{U|X}$ is sufficient for $P_{V|X}$. This concludes the proof of the Proposition 2. ∎
@@
$\inf\limits_{\rho_{U}}\sup\limits_{\{A_{y}\}}\,\mathbb{E}\left[\max\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,X),\rho_{U}\big{)}\right]\leqslant\varepsilon.$ (24)
@@
$\sum\limits_{y\in\mathsf{Y}}|\Gamma\big{(}(y,x),\rho_{U}\big{)}|.$
The above quantity is the induced $\ell_{1}$-norm distance between $\int_{\mathcal{A}_{y}}P_{V|X}(\mathsf{d}v|x)$ and $\int_{\mathsf{U}}\rho_{U}(y|u)P_{U|X}(\mathsf{d}u|x)$. Since, for all $x\in\mathsf{X}$,
@@
$\sum\limits_{y\in\mathsf{Y}}|\Gamma\big{(}(y,x),\rho_{U}\big{)}|=2\sum\limits_{y\in\mathsf{Y}:\Gamma((y,x),\rho_{U})\geqslant 0}\Gamma\big{(}(y,x),\rho_{U}\big{)},\quad\textrm{for all $x\in\mathsf{X}$}$
@@
$\sup\limits_{\{A_{y}\}}\mathbb{E}\sum\limits_{y\in\mathsf{Y}}|\Gamma\big{(}(y,X),\rho_{U}\big{)}|\leqslant 2|\mathsf{Y}|\sup\limits_{\{A_{y}\}}\mathbb{E}\left[\max\limits_{y\in\mathsf{Y}}\Gamma\big{(}(y,X),\rho_{U}\big{)}\right]\leqslant 2|\mathsf{Y}|\varepsilon,\quad\textrm{for all $x\in\mathsf{X}$}$